# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v2

@@ 3.1 Convergence of ChFSI with inexact subspace construction
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)\qquad\forall i=0,\dots,\infty$ (6)
We note that if $\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,$ remains nearly constant with iteration $i$, the right-hand side of the inequality in Eq. 6 keeps increasing as we approach the exact eigenspace and beyond a certain angle $\angle(\mathcal{S}^{(i)},\mathcal{S})$ this inequality gets violated. Consequently, the angle stops decreasing and we cannot approach the exact eigenspace beyond that point. To this end, for robust convergence, we require that $\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,$ also decreases as we approach the exact eigenspace, and we demonstrate that our proposed residual-based reformulation of Chebyshev filtered subspace iteration procedure (R-ChFSI) algorithm described subsequently accomplishes this. We further note that $\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)\geq 2\sqrt{2}$ and consequently for convergence on any approximation employed, we obtain the following necessary condition on $\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,$ from Eq. 6,
$\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,<\frac{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}{2\sqrt{2}}.$ (7)
@@ 3.2 Standard Eigenvalue Problems
3.2.1 Traditional Chebyshev filtering approach employing low-precision matrix-products
@@
$\underline{\textbf{Y}}_{k+1}^{(i)}=a_{k}{\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}+b_{k}\underline{\textbf{Y}}_{k}^{(i)}+c_{k}\underline{\textbf{Y}}_{k-1}^{(i)}$ (8)
@@
and with the initial conditions $\underline{\textbf{Y}}_{0}^{(i)}=\textbf{X}^{(i)}$ and $\underline{\textbf{Y}}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{H}-c\textbf{I})\textbf{X}^{(i)}$. This allows us to write a recurrence relation for $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
$\bm{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\bm{\Delta}_{k}^{(i)}+b_{k}\bm{\Delta}_{k}^{(i)}+c_{k}\bm{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}\right)$ (9)
with the initial conditions $\bm{\Delta}_{0}^{(i)}=\bm{\Delta}_{1}^{(i)}=0$. We now state some useful results required for further analysis
@@
Proof.
This is a straightforward application of the results derived in [69, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [69, section 3.6]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines. ∎
@@
The spectral norm of the error $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using lower precision arithmetic in matrix-products of the recurrence relation Eq. 3 satisfies $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ is some finite constant that depends on $k$
Proof.
We argue by induction that $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}$ where $\eta_{k}$ is some finite constant that depends on $k$. Recall from the initial conditions on Eq. 9, we have $\bm{\Delta}_{0}^{(i)}=\bm{\Delta}_{1}^{(i)}=0$. Further, we assume the induction hypothesis for $k=r-1$ and $k=r$ i.e., the inequalities $\lVert\bm{\Delta}_{r-1}^{(i)}\rVert\,\leq\gamma_{m}\eta_{r-1}$ and $\lVert\bm{\Delta}_{r}^{(i)}\rVert\,\leq\gamma_{m}\eta_{r}$ are true and now we need to show $\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,\leq\gamma_{m}\eta_{r+1}$. Subsequently, from the induction hypothesis, we have the following expression for $\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ using Eq. 9
$\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ $\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\eta_{r}\gamma_{m}+\left\lvert b_{r}\right\rvert\,\eta_{r}\gamma_{m}+\lvert c_{r}\rvert\,\eta_{r-1}\gamma_{m}+\left\lvert a_{r}\right\rvert\,\lVert{\textbf{H}\otimes\underline{\textbf{Y}}_{r}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{r}^{(i)}\rVert\,$ (10)
$\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\eta_{r}\gamma_{m}+\left\lvert b_{r}\right\rvert\,\eta_{r}\gamma_{m}+\lvert c_{r}\rvert\,\eta_{r-1}\gamma_{m}+\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\lVert\underline{\textbf{Y}}_{r}^{(i)}\rVert\,\gamma_{m}$ (11)
@@
where we have used Lemma 3.1. We further note that $\lVert\underline{\textbf{Y}}_{r}^{(i)}\rVert\,=\lVert\bm{\Delta}_{r}^{(i)}+\textbf{Y}_{r}^{(i)}\rVert\,\leq\gamma_{m}\eta_{r}+\lVert\textbf{Y}_{r}^{(i)}\rVert\,\leq\gamma_{m}\eta_{r}+\lVert C_{r}(\textbf{H})\textbf{X}^{(i)}\rVert\,\leq\gamma_{m}\eta_{r}+\lVert C_{r}(\textbf{H})\rVert\,$ and consequently we can write
$\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ $\leq\left[\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\eta_{r}+\left\lvert b_{r}\right\rvert\,\eta_{r}+\lvert c_{r}\rvert\,\eta_{r-1}+\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,(\gamma_{m}\eta_{r}+\lVert C_{r}(\textbf{H})\rVert\,)\right]\gamma_{m}$ (12)
Defining $\eta_{r+1}=\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\eta_{r}+\left\lvert b_{r}\right\rvert\,\eta_{r}+\lvert c_{r}\rvert\,\eta_{r-1}+\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,(\gamma_{m}\eta_{r}+\lVert C_{r}(\textbf{H})\rVert\,)$, we have $\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,\leq\gamma_{m}\eta_{r+1}$ from Eq. 12. Hence using induction, we argue that $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}$ for $k=0,1,\dots,p$
@@
We now discuss the implications of the above Theorem 3.2 within the context of Eq. 6. Without loss of generality, we assume that our initial guess of trial subspace $\mathcal{S}^{(i)}$ at $i=0$ satisfies the inequality below as stated in Eq. 6.
@@
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (13)
@@
$\angle(\mathcal{S}^{(i)},\mathcal{S})\approx\frac{\gamma_{m}\eta_{p}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}$ (14)
@@ 3.2.2 Proposed residual-based Chebyshev filtering approach
$\textbf{R}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\bm{\Lambda}^{(i)})=\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)}\;\;\;\text{for}\;\;\;k=0,\dots,p$ (15)
where we have defined $\bm{\Lambda}_{k}^{(i)}=C_{k}(\bm{\Lambda}^{(i)})$ and recall $p$ is the maximum Chebyshev polynomial degree used in the subspace construction step.
@@
The recurrence relation given by Eq. 3 can be reformulated in terms of the residuals defined by $\textbf{R}_{k}^{(i)}=\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)}$ as
$\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{H}\textbf{R}_{k}^{(i)}+b_{k}\textbf{R}_{k}^{(i)}+c_{k}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\bm{\Lambda}_{k}^{(i)}$ (16)
$\bm{\Lambda}_{k+1}^{(i)}$ $=a_{k}\bm{\Lambda}_{k}^{(i)}\bm{\Lambda}^{(i)}+b_{k}\bm{\Lambda}_{k}^{(i)}+c_{k}\bm{\Lambda}_{k-1}^{(i)}$ (17)
where $\textbf{R}_{0}^{(i)}=0$ and $\textbf{R}_{1}^{(i)}=\frac{\sigma_{1}}{e}\textbf{R}^{(i)}$ with $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}^{(i)}$ and further we have $a_{k}=2\sigma_{k+1}/e$, $b_{k}=-2\sigma_{k+1}c/e$, and $c_{k}=-\sigma_{k}\sigma_{k+1}$
Proof.
The recurrence relation for $\bm{\Lambda}_{k}^{(i)}=C_{k}(\bm{\Lambda}^{(i)})$ can be written as
$\bm{\Lambda}_{k+1}^{(i)}=a_{k}\bm{\Lambda}_{k}^{(i)}\bm{\Lambda}^{(i)}+b_{k}\bm{\Lambda}_{k}^{(i)}+c_{k}\bm{\Lambda}_{k-1}^{(i)}$
@@
$\textbf{Y}_{k+1}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k+1}^{(i)}$ $=a_{k}(\textbf{H}\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)}\bm{\Lambda}^{(i)})+b_{k}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)})+c_{k}(\textbf{Y}_{k-1}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)})$ (18)
$\implies\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{H}\textbf{R}_{k}^{(i)}+b_{k}\textbf{R}_{k}^{(i)}+c_{k}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\bm{\Lambda}_{k}^{(i)}$ (19)
After computing $\textbf{R}_{p}^{(i)}$ using this recurrence relation we can now recover the desired subspace $\textbf{Y}_{p}^{(i)}$ using the relation in Eq. 15 i.e., $\textbf{Y}_{p}^{(i)}=\textbf{R}_{p}^{(i)}+\textbf{X}^{(i)}\bm{\Lambda}_{p}^{(i)}$. We note that using approximate matrix multiplication for evaluating $\textbf{H}\textbf{R}_{k}^{(i)}$ we can write the following recurrence relation
$\underline{\textbf{R}}_{k+1}^{(i)}=a_{k}{\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}+b_{k}\underline{\textbf{R}}_{k}^{(i)}+c_{k}\underline{\textbf{R}}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\bm{\Lambda}_{k}^{(i)}$ (20)
and consequently, we have $\underline{\textbf{Y}}_{k}^{(i)}=\underline{\textbf{R}}_{k}^{(i)}+\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)}$. This allows us to write a recurrence relation for $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}=\underline{\textbf{R}}_{k}^{(i)}-\textbf{R}_{k}^{(i)}$ as
$\bm{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\bm{\Delta}_{k}^{(i)}+b_{k}\bm{\Delta}_{k}^{(i)}+c_{k}\bm{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{R}}_{k}^{(i)}\right)$ (21)
where $\bm{\Delta}_{0}^{(i)}=\bm{\Delta}_{1}^{(i)}=0$. We now prove a result useful for further analysis

### v3

@@ 3.1 Convergence of ChFSI with inexact subspace construction
(5) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,\!>\!\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)\;\forall i=0,\dots,\infty$
We note that if $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ remains nearly constant with iteration $i$, the right-hand side of the inequality in Eq. 5 keeps increasing as we approach the exact eigenspace and beyond a certain angle $\angle(\mathcal{S}^{(i)},\mathcal{S})$ this inequality gets violated. Consequently, the angle stops decreasing and we cannot approach the exact eigenspace beyond that point. To this end, for robust convergence, we require that $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ also decreases as we approach the exact eigenspace, and we demonstrate that our proposed residual-based reformulation of Chebyshev filtered subspace iteration procedure (R-ChFSI) algorithm described subsequently accomplishes this. We further note that $\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)\geq 2\sqrt{2}$ and consequently for convergence with approximations, we obtain the following necessary condition on $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ from Eq. 5,
(6) $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,<\frac{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}{2\sqrt{2}}.$
@@ 3.2 Standard Eigenvalue Problems
3.2.1 Traditional Chebyshev filtering method employing low-precision matrix-products
@@
(7) $\underline{\textbf{Y}}_{k+1}^{(i)}=a_{k}{\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}+b_{k}\underline{\textbf{Y}}_{k}^{(i)}+c_{k}\underline{\textbf{Y}}_{k-1}^{(i)}$
@@
and with the initial conditions $\underline{\textbf{Y}}_{0}^{(i)}=\textbf{X}^{(i)}$ and $\underline{\textbf{Y}}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{H}-c\textbf{I})\textbf{X}^{(i)}$.
We now state some useful results required for further analysis
@@
Proof 3.4.
This is a straightforward application of the results derived in [14, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [14, section 3.6]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines.
Lemma 3.5.
Consider a recurrence relation of the form
(8) $\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\textbf{e}_{k}^{(i)}$
with $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq h_{0}\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+h_{1}^{(i)}\;\forall k=1,2,\dots,p$. We can then write
(9) $\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\right)+g_{k}^{1}h_{1}^{(i)}$
Proof 3.6.
From Eq. 8 we can write
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,$ $\leq\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lvert b_{k}\rvert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lvert c_{k}\rvert\,\lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,+\lvert a_{k}\rvert\,\lVert\textbf{e}_{k}^{(i)}\rVert\,$
$\leq(\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,+\lvert b_{k}\rvert\,+h_{0})\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lvert c_{k}\rvert\,\lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,+\lvert a_{k}\rvert\,h_{1}^{(i)}$
We define the following
$\textbf{d}_{k}^{(i)}$ $=\begin{bmatrix}\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ \lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,\end{bmatrix}$ $\textbf{F}_{k}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,+\lvert b_{k}\rvert\,+\lvert a_{k}\rvert\,h_{0}&\lvert c_{k}\rvert\,\\ 1&0\end{bmatrix}$ $\textbf{E}_{k}^{(i)}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,h_{1}^{(i)}\\ 0\end{bmatrix}$
we can now write
$\lVert\textbf{d}_{k+1}^{(i)}\rVert\,$ $\leq\lVert\textbf{F}_{k}\textbf{d}_{k}^{(i)}+\textbf{E}_{k}^{(i)}\rVert\,\leq\lVert\textbf{F}_{k}\rVert\,\lVert\textbf{d}_{k}^{(i)}\rVert\,+\lvert a_{k}\rvert\,h_{1}^{(i)}$
Consequently, we now have
$\lVert\textbf{d}_{k+1}^{(i)}\rVert\,$ $\leq\left(\prod_{j=1}^{k}\lVert\textbf{F}_{j}\rVert\,\right)\lVert\textbf{d}_{1}^{(i)}\rVert\,+\sum_{j=1}^{k}\left(\prod_{r=j+1}^{k}\lVert\textbf{F}_{r}\rVert\,\right)\lvert a_{j}\rvert\,h_{1}^{(i)}$
We note that $\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq\lVert\textbf{d}_{k+1}^{(i)}\rVert\,$ and $\lVert\textbf{d}_{1}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,$. This results in
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq\left(\prod_{j=1}^{k}\lVert\textbf{F}_{j}\rVert\,\right)\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\right)+\sum_{j=1}^{k}\left(\prod_{r=j+1}^{k}\lVert\textbf{F}_{r}\rVert\,\right)\lvert a_{j}\rvert\,h_{1}^{(i)}$
Defining $g_{k}^{0}=\prod_{j=1}^{k}\lVert\textbf{F}_{j}\rVert\,$ and $g_{k}^{1}=\sum_{j=1}^{k}\prod_{r=j+1}^{k}\lVert\textbf{F}_{r}\rVert\,\lvert a_{j}\rvert\,$ we have
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\right)+g_{k}^{1}h_{1}^{(i)}$
@@
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using lower precision arithmetic in matrix-products of the recurrence relation Eq. 2 satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ is some finite constant that depends on $k$
Proof 3.8.
Using Eq. 2 and Eq. 7 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
(10) $\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}\right)$
@@
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=\boldsymbol{\Delta}_{1}^{(i)}=0$. We note that this recurrence relation is of the same form as that of Eq. 8 in Lemma 3.5 with $\textbf{e}_{k}^{(i)}={\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}$. From Lemma 3.3 we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,=\lVert{\textbf{H}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,$. We further note that
$\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{k}^{(i)}+\textbf{Y}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert\textbf{Y}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert C_{k}(\textbf{H})\textbf{X}^{(i)}\rVert\,\\ \leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert C_{k}(\textbf{H})\rVert\,$
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\gamma_{m}\lVert\textbf{H}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$. Using Lemma 3.5 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,=0$, $h_{0}=\gamma_{m}\lVert\textbf{H}\rVert\,$ and $h_{1}^{(i)}=\gamma_{m}\lVert\textbf{H}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}$ with $\eta_{k}=\lVert\textbf{H}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}$.
@@
We now discuss the implications of the Theorem 3.7 within the context of Eq. 5. Without loss of generality, we assume that our initial guess of trial subspace $\mathcal{S}^{(i)}$ at $i=0$ satisfies the inequality below as stated in Eq. 5.
@@
(11) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$
@@
(12) $\angle(\mathcal{S}^{(i)},\mathcal{S})\approx\frac{\gamma_{m}\eta_{p}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}$
@@ 3.2.2 Proposed residual-based Chebyshev filtering approach
(13) $\textbf{R}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\boldsymbol{\Lambda}^{(i)})=\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}\;\;\;\text{for}\;\;\;k=0,\dots,p$
where we have defined $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ and recall $p$ is the maximum Chebyshev polynomial degree used in the subspace construction step.
@@
The recurrence relation given by Eq. 2 can be reformulated in terms of the residuals defined by $\textbf{R}_{k}^{(i)}=\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$ as
(14) $\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{H}\textbf{R}_{k}^{(i)}+b_{k}\textbf{R}_{k}^{(i)}+c_{k}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
(15) $\boldsymbol{\Lambda}_{k+1}^{(i)}$ $=a_{k}\boldsymbol{\Lambda}_{k}^{(i)}\boldsymbol{\Lambda}^{(i)}+b_{k}\boldsymbol{\Lambda}_{k}^{(i)}+c_{k}\boldsymbol{\Lambda}_{k-1}^{(i)}$
where $\textbf{R}_{0}^{(i)}=0$ and $\textbf{R}_{1}^{(i)}=\frac{\sigma_{1}}{e}\textbf{R}^{(i)}$ with $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$ and further we have $a_{k}=2\sigma_{k+1}/e$, $b_{k}=-2\sigma_{k+1}c/e$, and $c_{k}=-\sigma_{k}\sigma_{k+1}$
Proof 3.10.
The recurrence relation for $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ can be written as
$\boldsymbol{\Lambda}_{k+1}^{(i)}=a_{k}\boldsymbol{\Lambda}_{k}^{(i)}\boldsymbol{\Lambda}^{(i)}+b_{k}\boldsymbol{\Lambda}_{k}^{(i)}+c_{k}\boldsymbol{\Lambda}_{k-1}^{(i)}$
@@
$\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{H}\textbf{R}_{k}^{(i)}+b_{k}\textbf{R}_{k}^{(i)}+c_{k}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
After computing $\textbf{R}_{p}^{(i)}$ using this recurrence relation we can now recover the desired subspace $\textbf{Y}_{p}^{(i)}$ using the relation in Eq. 13 i.e., $\textbf{Y}_{p}^{(i)}=\textbf{R}_{p}^{(i)}+\textbf{X}^{(i)}\boldsymbol{\Lambda}_{p}^{(i)}$. We note that using approximate matrix multiplication for evaluating $\textbf{H}\textbf{R}_{k}^{(i)}$ we can write the following recurrence relation
(16) $\underline{\textbf{R}}_{k+1}^{(i)}=a_{k}{\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}+b_{k}\underline{\textbf{R}}_{k}^{(i)}+c_{k}\underline{\textbf{R}}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
and consequently, we have $\underline{\textbf{Y}}_{p}^{(i)}=\underline{\textbf{R}}_{p}^{(i)}+\textbf{X}^{(i)}\boldsymbol{\Lambda}_{p}^{(i)}$.
We now prove a result useful for further analysis